OpenStudy (anonymous):

completing the square. x^2+6x+9=1 x= and x= help please i will give medal

OpenStudy (raden):

hint : x^2+6x+9 = (x+3)^2

OpenStudy (anonymous):

so x+3=1 and x-3=1 so ssolve or x right

OpenStudy (raden):

x+3=1 or x+3=-1 solve for x

OpenStudy (anonymous):

-2 and -4

OpenStudy (raden):

yes, you are right

OpenStudy (anonymous):

i got the question wrong that isnt how you complete the square

OpenStudy (anonymous):

:( noone knows how to do it

OpenStudy (raden):

x^2+6x+9=1 actually, the terms in left side already to completing the square if u expand the term of (x+3)^2 , it is x^2+6x+9 :)

OpenStudy (anonymous):

so simplify to x= and x=

OpenStudy (raden):

from (x+3)^2 = 1 take the square root , it can be x+3 = +- sqrt(1)

OpenStudy (anonymous):

huh